# Physics-informed graph neural networks for flow field estimation in carotid arteries

### v1

@@ Abstract
{IEEEkeywords}
4D flow MRI, geometric deep learning, physics-informed machine learning, carotid arteries.
Figure 1: Overview. We represent the carotid artery by $n$ un-ordered points with $c_{\text{in}}$-dimensional input features, one of which is patient-specific inflow $u_{\text{in}}$ (shown in green). PointNet++ learns to map the input features to 3D velocity vectors using hierarchical down-sampling (pooling) and up-sampling (interpolation) layers. PointNet++ is comprised of learnable functions which we choose as vanilla multilayer perceptrons (MLP) or $\mathrm{E}(3)$-steerable [1] MLPs. We train PointNet++ with ground truth velocity fields obtained in-vivo via 4D flow MRI. Based on the Navier-Stokes equations for incompressible fluids that govern arterial blood flow, we regularise the training loss with residuals using discretised differential operators.
@@ 1 Introduction
Cardiovascular disease is the leading cause of death worldwide. In many cases, it is characterised by an accumulation of plaque in the arterial wall, which results in narrowing of the blood vessel. In extreme cases, such stenosis causes a shortage of oxygenated blood supply to downstream organs which can lead to myocardial infarction, peripherial artery disease, or ischemic stroke [2]. Ischemic stroke is often the result of atherosclerosis in the carotid arteries. Carotid artery geometry has been identified as an indicator for atherosclerosis [3] and disturbed blood flow [4]. Arterial blood flow, together with derived quantities like wall shear stress and oscillatory shear index, has been shown to correlate with initiation and progression of atherosclerosis [5]. Thus, insight into patient-specific blood flow is invaluable for diagnosis, treatment and prevention of cardiovascular disease.
@@
Information about patient-specific hemodynamics can be clinically obtained in-vivo, via invasive measurements such as pressure sensors in catheterisation or via non-invasive imaging such as Doppler ultrasound, particle image velocimetry [6] or 4D flow magnetic resonance imaging (MRI) [2]. In particular, 4D flow MRI allows for the quantification of directional blood flow and has emerged as a leading technique for in-vivo hemodynamics measurement [7]. However, 4D flow requires expensive, specialised software and expert knowledge, in particular for the setting of scan parameters. More importantly, the number of patients for whom hemodynamics measurements could potentially provide additional diagnostic information commonly exceeds a hospital’s capacity for performing 4D flow MRI, making the number of available scanners a bottleneck. Furthermore, 4D flow MRI is prone to measurement noise and inaccurate electrocardiogram (ECG) gating [7].
@@
Alternatively, information about patient-specific hemodynamics can be obtained in-silico via computational fluid dynamics (CFD) based on anatomical computed tomography (CT) [8] or MRI [9]. CFD simulation based on 3D vascular models extracted from anatomical imaging is a powerful tool to estimate hemodynamic quantities in-silico. CFD allows for accurate, physics-conforming blood flow estimation. Its downside are long runtimes and high computational demand, as well as sensitivity to modelling choices, such as discretisation and boundary conditions, making it difficult to compare results across practitioners. Indeed, it has been shown that there is high variability across CFD simulations performed by different research groups [10]. This underlines the fact that while being a powerful model for blood flow estimation, CFD is not comparable to in-vivo measurement and must be validated via reference measurements, like 4D flow MRI. These limitations can be prohibitive for widespread use in clinical practice.
@@
In this work, we aim to address the individual shortcomings of 4D flow MRI and CFD by creating an in-silico surrogate model for hemodynamic flow field estimation. Based on the observation that hemodynamics largely depend on artery shape and flow conditions, we propose a machine learning model that estimates patient-specific blood flow from personalised 3D vascular models. We demonstrate this approach in the carotid artery and train our machine learning model on 4D flow measurements in the carotid bifurcation of 234 subjects. We employ a powerful class of neural networks based on the PointNet++ architecture, which has been shown to excel at dealing with large point clouds [11, 12]. Similarly as in CFD simulation [13], our machine-learning-based flow predictions are conditioned on patient-specific boundary conditions. Once trained, our model can quickly produce flow field estimates for vascular anatomy based on boundary conditions that can be cheaply obtained, e.g., via the widely available Doppler ultrasound. An overview of our method is provided in Fig. 1.
@@
Learning a relation between vascular geometry and hemodynamics from in-vivo data has the clear advantage that the estimated flow fields, unlike those obtained from CFD, do not depend on geometric and physical modelling choices. However, it poses the practical challenges that (1) training data is limited because 4D flow MRI is not performed on large cohorts of subjects and (2) measurement noise results in data pollution which can make training challenging. To address the former, we incorporate group equivariance by embedding steerable $\mathrm{E}(3)$-equivariant layers [1] in the PointNet++ architecture, which has been shown to facilitate extracting knowledge out of small hemodynamic datasets [14]. To address the latter, we include Navier-Stokes-based loss regularisation to mitigate the effect of un-physical noise in the training data. We perform experiments using conservation of mass as training loss regularisation alongside conservation of momentum where we find it to be independently beneficial for accuracy. We achieve this via discretisation of differential operators for which we derive a novel discretisation scheme using computationally efficient neighbourhood queries. The resulting model is fast, patient-specific and produces physics-conforming estimates.
@@ 2 Related works
Recently, considerable attention has been placed on machine learning methods for hemodynamics estimation. These approaches can be divided into two categories: (1) transductive instance optimisation with physics-informed regularisation (similar in scope to CFD) and (2) inductive, generalising feed-forward methods that learn a relation between anatomy and hemodynamics from training data. Methods in the first category can incorporate personalised boundary conditions but lack the ability to generalise and require re-training for each case. In such, they can be useful in settings where the hemodynamic parameters are partially known [15, 16, 17]. In contrast, methods in the latter category learn to generalise to unseen data and aim to create fast, compute-efficient surrogate models which may replace compute-intensive CFD.
In this work we focus on the second category: learning a relation between subjects’ vascular geometry and hemodynamics using neural networks. These neural networks operate on 3D representations of the vessels and estimate blood-flow-related quantities on the vessel wall or flow fields in the interior of the vessel. Besides efforts to estimate surface quantities like wall shear stress [18] and endothelial cell activation potential [12], previous works have focussed on the estimation of volumetric (vector) fields. Among these, Liang et al. [19] and Wang et al. [20] trained fully-connected neural networks operating on 3D point-cloud representations of the carotid artery and thoracic aorta, respectively, to estimate pressure and velocity fields. Maul et al. [21] employed an octree-based neural network followed by trilinear interpolation to learn a continuous solution operator for the Navier-Stokes equations in synthetic vascular trees. Li et al. [22] used a PointNet-like [23] architecture to estimate pressure and velocity fields in the coronary artery and synthetic cerebral aneurysm, respectively. Zhang et al. [11] composed PointNet++ [24] and a Navier-Stokes physics-informed neural network (PINN) which learned to estimate the velocity field in 3D models of the abdominal aorta. In an earlier study [14], we used a multiscale steerable $\mathrm{E}(3)$-equivariant graph neural network (SEGNN) [1] to estimate velocity fields in synthetic coronary arteries.
@@ 3 Materials and method
We included data of 234 subjects. All subjects participated in a long-term follow-up study evaluating cardiovascular risk in patients with familial hypercholesterolemia (FH) in whom statin treatment was initiated in childhood and in their unaffected siblings [25]. All procedures were approved by the local institutional review board (METC) of the Amsterdam University Medical Center and were carried out according to the declaration of Helsinki. In all subjects, 4D flow MRI and 3D black blood MRI volumes were acquired.
@@
4D flow scans were acquired using the following parameters: TR = 7.8 ms, TE = 4.6 ms, flip angle = 8°, VENC = 150 $\frac{\text{cm}}{\text{s}}$, $0.8\times 0.8\times 0.8$ $\text{mm}^{3}$ spatial and 80 ms temporal resolution. Retrospective ECG-triggering was used for cardiac synchronisation. Phase-offset corrections were performed during reconstruction. The 4D flow scan was accelerated using a k-t undersampling scheme of factor R = 8 and a k-t PCA reconstruction [26]. Images were reconstructed using CRecon (Gyrotools, Zurich, Switzerland) with a regularisation factor of $r=0.1$. The total scan time was 10 minutes. An nnU-Net model [27] was used to automatically segment a region-of-interest from an estimated 3 cm below to 2 cm above the carotid bifurcation. Triangular meshes for the carotid arteries were obtained from this segmentation and used to mask the velocity vector data. We chose the peak systolic time-frame as target in this study, which was defined as the time-frame with the highest spatially averaged velocity.
@@
Black blood MRI scans were performed using the following parameters: TR = 10 ms, TE = 3.4 ms, flip angle = 6° and 0.5 mm $\times$ 0.5 mm $\times$ 0.5 mm spatial resolution. The k-space was under-sampled R = 5 times with a Poisson disk pattern and reconstructed using a compressed sensing algorithm for which the previously developed PROUD scanner software patch was used [28, 29]. Reconstruction was carried out in MATLAB, using MRecon (Gyrotools, Zurich, Switzerland) and the BART reconstruction toolbox. In all 3D black blood MRI volumes, the internal carotid artery (ICA), external carotid artery (ECA) and common carotid artery (CCA) were automatically segmented using a previously developed deep learning-based algorithm [30]. Meshes were subsequently cut at 3 cm below and 2 cm above the carotid bifurcation to match the region-of-interest obtained from 4D flow MRI.
@@
After visual inspection of the meshes obtained from 4D flow MRI we excluded 162 arteries for the current study. Reasons were incompleteness of the triangular mesh, inclusion of arteries other than the carotid artery and meshing errors beyond repair, that would affect downstream analysis. Hence, the final dataset consisted of 128 left and 154 right carotid arteries, for which paired meshes were available from 4D flow MRI and black blood MRI. For both 4D flow and black blood MRI, we obtained centerline points by binning and averaging geodesic segments of the artery surface, computed via the vector heat method [31]. Connecting these centerline points by line elements allows us to query the distance to the centerline and thus vessel radius using trigonometry.
@@ 3.2 Machine learning model
In the following section we introduce our machine learning model for learning to predict velocity fields in 3D artery models. We propose a neural network that maps a collection of input features, defined on a set of points, which describe (local) artery geometry within the artery lumen, to a collection of output features representing 3D velocity vectors of blood flow. Note that neither rotation nor translation of the input geometry should influence the relative structure of the predicted vector field. We address this by a steerable, $\mathrm{SE}(3)$-equivariant neural network that processes descriptive, geometric features independent of orientation in ambient space. Furthermore, we include the Navier-Stokes equations as loss regularisation.
@@ 3.2.1 Input features
We describe arterial geometry locally by point-wise feature vectors $f^{p}\in\mathbb{R}^{c}$ for each point $p\in\mathcal{P}_{0}$, where $c$ is the channel size and $\mathcal{P}_{0}$ a finite set of points within the artery lumen. We describe each point by its relative position (compare [14]) to the artery inlet, lumen wall, outlets and centerline. We additionally encode relative position to the ICA outlet, ECA outlet and append a field of zeros (or ones) if the artery is a left (or right) carotid artery (respectively). We do so to account for possible physiological differences with respect to upstream or downstream vasculature and their effect on the flow field. All these features represent relative positions and are thus invariant under translation of the 3D geometry.
@@
Furthermore, we include boundary condition features, as constant scalar fields extended over all points, namely mean velocity and standard deviation over the artery inlet. The latter is to simulate velocity measurements in an axial slice of the artery. Such measurements are commonly used in CFD to get subject-specific flow boundary conditions, often via ultrasound [13]. Conditioning on different inflow velocities is important because boundary conditions vary between subjects and greatly influence blood flow.
@@ 3.2.2 Architecture
We propose an encoder-decoder model (following PointNet++ [24]) with a contracting and expanding pathway. Given an un-ordered set $\mathcal{P}_{0}=\{p^{1},\dots,p^{n}\}$ of point coordinates $p=(p_{1},p_{2},p_{3})^{\mathsf{T}}\in\mathbb{R}^{3}$ within a subject’s artery lumen, we build a nested hierarchy of $h$ sub-sampled sets
@@ 3.2.3 Steerable equivariant PointNet++
We introduce an equivariant extension of PointNet++ using steerable representations. Generally, a function is called equivariant to a symmetry group, if group actions applied to the inputs result in the equivalent transformation of the outputs. By choosing $\phi$, $\psi$ as steerable MLPs we, in turn, render our PointNet++ $\mathrm{O}(3)$-equivariant. Steerable MLPs that are equivariant under transformations of the orthogonal group $\mathrm{O}(3)$ (rotations and reflections) can be constructed by interleaving Clebsch-Gordan tensor products, gated activation functions and batch normalisation [1]. They require expressing feature vectors throughout the network as steerable tensors (e.g. collections of scalars $s$ and 3D vectors $v$)

### v2

@@ Abstract
keywords:
[…]
@@ 1 Introduction
Cardiovascular disease is the leading cause of death worldwide. In many cases, it is characterised by an accumulation of plaque in the arterial wall, which results in narrowing of the blood vessel. In extreme cases, such stenosis causes a shortage of oxygenated blood supply to downstream organs which can lead to myocardial infarction, peripherial artery disease or ischemic stroke Saba et al. (2023). Ischemic stroke is often the result of atherosclerosis in the carotid arteries. Carotid artery geometry has been identified as an indicator for atherosclerosis Strecker et al. (2020) and disturbed blood flow Lee et al. (2008). Arterial blood flow, together with derived quantities like wall shear stress and oscillatory shear index, has been shown to correlate with initiation and progression of atherosclerosis Cibis et al. (2016). Thus, insight into patient-specific blood flow is invaluable for diagnosis, treatment and prevention of cardiovascular disease.
@@
Information about patient-specific hemodynamics can be clinically obtained in-vivo, via invasive measurements such as pressure sensors in catheterisation or via non-invasive imaging such as Doppler ultrasound, particle image velocimetry Engelhard et al. (2018) or 4D flow magnetic resonance imaging (MRI) Saba et al. (2023). In particular, 4D flow MRI allows for the quantification of directional blood flow and has emerged as a leading technique for in-vivo hemodynamics measurement Bissell et al. (2023). However, 4D flow requires expensive, specialised software and expert knowledge, in particular for the setting of scan parameters. More importantly, the number of patients for whom hemodynamics measurements could potentially provide additional diagnostic information commonly exceeds a hospital’s capacity for performing 4D flow MRI, making the number of available scanners a bottleneck. Furthermore, 4D flow MRI is prone to measurement noise and inaccurate electrocardiogram (ECG) gating Bissell et al. (2023).
@@
Alternatively, information about patient-specific hemodynamics can be obtained in-silico via computational fluid dynamics (CFD) based on anatomical computed tomography (CT) Li et al. (2024) or MRI Taylor et al. (2023). CFD simulation based on 3D vascular models extracted from anatomical imaging is a powerful tool to estimate hemodynamic quantities in-silico. CFD allows for accurate, physics-conforming blood flow estimation. Its downside are long runtimes and high computational demand, as well as sensitivity to modelling choices, such as discretisation and boundary conditions, making it difficult to compare results across practitioners. Indeed, it has been shown that there is high variability across CFD simulations performed by different research groups Valen-Sendstad et al. (2018). This underlines the fact that while being a powerful model for blood flow estimation, CFD is not comparable to in-vivo measurement and must be validated via reference measurements, like 4D flow MRI. These limitations can be prohibitive for widespread use in clinical practice.
Figure 1: Overview. We represent the carotid artery by $n$ un-ordered points with $c_{\text{in}}$-dimensional input features, one of which is patient-specific inflow $u_{\text{in}}$ (shown in green). PointNet++ learns to map the input features to 3D velocity vectors using hierarchical down-sampling (pooling) and up-sampling (interpolation) layers. PointNet++ is comprised of learnable functions which we choose as vanilla multilayer perceptrons (MLP) or steerable Brandstetter et al. (2022) MLPs. We train PointNet++ with ground truth velocity fields obtained in-vivo via 4D flow MRI. Based on the Navier-Stokes equations for incompressible fluids that govern arterial blood flow, we regularise the training loss with residuals using discretised differential operators.
@@
In this work, we aim to address the individual shortcomings of 4D flow MRI and CFD by creating an in-silico surrogate model for hemodynamic flow field estimation. Based on the observation that hemodynamics largely depend on artery shape and flow conditions, we propose a machine learning model that estimates patient-specific blood flow from personalised 3D vascular models. We demonstrate this approach in the carotid artery and train our machine learning model on 4D flow measurements in the carotid bifurcation of 234 subjects. We employ a powerful class of neural networks based on the PointNet++ architecture, which has been shown to excel at dealing with large point clouds Zhang et al. (2023); Morales Ferez et al. (2021). Notably, the model processes data in the form of point clouds rather than meshes to avoid overfitting to vertex connectivity and to overcome restrictions in receptive field. Similarly as in CFD simulation Lopes et al. (2020), our machine-learning-based flow predictions are conditioned on patient-specific boundary conditions. Once trained, our model can quickly produce flow field estimates for vascular anatomy based on boundary conditions that can be cheaply obtained, e.g., via the widely available Doppler ultrasound. An overview of our method is provided in Fig. 1.
@@
Learning a relation between vascular geometry and hemodynamics from in-vivo data has the clear advantage that the estimated flow fields, unlike those obtained from CFD, do not depend on geometric and physical modelling choices. However, it poses the practical challenges that (1) training data is limited because 4D flow MRI is not performed on large cohorts of subjects and (2) measurement noise results in data pollution which can make training challenging. To address the former, we incorporate group equivariance by embedding steerable $\mathrm{E}(3)$-equivariant layers Brandstetter et al. (2022) in the PointNet++ architecture, which has been shown to facilitate extracting knowledge out of small hemodynamic datasets Suk et al. (2023). To address the latter, we include Navier-Stokes-based loss regularisation to mitigate the effect of un-physical noise in the training data. We perform experiments using conservation of mass as training loss regularisation alongside conservation of momentum. We achieve this via discretisation of differential operators for which we derive a novel discretisation scheme using computationally efficient neighbourhood queries. The resulting model is fast, patient-specific and produces physics-conforming estimates.
@@ 2 Related works
Recently, considerable attention has been placed on machine learning methods for hemodynamics estimation. These approaches can be divided into two categories: (1) transductive instance optimisation with physics-informed regularisation (similar in scope to CFD) and (2) inductive, generalising feed-forward methods that learn a relation between anatomy and hemodynamics from training data. Methods in the first category can incorporate personalised boundary conditions but lack the ability to generalise and require re-training for each case. In such, they can be useful in settings where the hemodynamic parameters are partially known Raissi et al. (2020); Fathi et al. (2020); Kontogiannis and Juniper (2022). In contrast, methods in the latter category learn to generalise to unseen data and aim to create fast, compute-efficient surrogate models which may replace compute-intensive CFD.
In this work we focus on the second category: learning a relation between subjects’ vascular geometry and hemodynamics using neural networks. These neural networks operate on 3D representations of the vessels and estimate blood-flow-related quantities on the vessel wall or flow fields in the interior of the vessel. Besides efforts to estimate surface quantities like wall shear stress Suk et al. (2024) and endothelial cell activation potential Morales Ferez et al. (2021), previous works have focussed on the estimation of volumetric (vector) fields. Among these, Liang et al. (2020) and Wang et al. (2023) trained fully-connected neural networks operating on 3D point-cloud representations of the carotid artery and thoracic aorta, respectively, to estimate pressure and velocity fields. Maul et al. (2023) employed an octree-based neural network followed by trilinear interpolation to learn a continuous solution operator for the Navier-Stokes equations in synthetic vascular trees. Li et al. (2021) used a PointNet-like Qi et al. (2016) architecture to estimate pressure and velocity fields in the coronary artery and synthetic cerebral aneurysm, respectively. Zhang et al. (2023) composed PointNet++ Qi et al. (2017) and a Navier-Stokes physics-informed neural network (PINN) which learned to estimate the velocity field in 3D models of the abdominal aorta. In an earlier study Suk et al. (2023), we used a multiscale steerable $\mathrm{E}(3)$-equivariant graph neural network (SEGNN) Brandstetter et al. (2022) to estimate velocity fields in synthetic coronary arteries.
@@ 3 Materials and methods
We selectively included data of 234 subjects. All subjects participated in a long-term follow-up study evaluating cardiovascular risk in patients with familial hypercholesterolemia (FH) in whom statin treatment was initiated in childhood and in their unaffected siblings Luirink et al. (2019). All procedures were approved by the local institutional review board (METC) of the Amsterdam University Medical Center and were carried out according to the declaration of Helsinki. In all subjects, 4D flow MRI and 3D black blood MRI volumes were acquired.
@@
4D flow scans were acquired using the following parameters: TR = 7.8 ms, TE = 4.6 ms, flip angle = 8°, VENC = 150 $\frac{\text{cm}}{\text{s}}$, $0.8\times 0.8\times 0.8$ $\text{mm}^{3}$ spatial and 80 ms temporal resolution. Retrospective ECG-triggering was used for cardiac synchronisation. Phase-offset corrections were performed during reconstruction. The 4D flow scan was accelerated using a k-t undersampling scheme of factor R = 8 and a k-t PCA reconstruction Pedersen et al. (2009). Images were reconstructed using CRecon (Gyrotools, Zurich, Switzerland) with a regularisation factor of $r=0.1$. The total scan time was 10 minutes. An nnU-Net model Isensee et al. (2020) was used to automatically segment a region-of-interest from an estimated 3 cm below to 2 cm above the carotid bifurcation. Triangular meshes for the carotid arteries were obtained from this segmentation and used to mask the velocity vector data. We chose the peak systolic time frame as target in this study, which was defined as the time frame with the highest spatially averaged velocity.
To facilitate visualisation of the flow fields, we created volumetric meshes from the surface meshes using TetGen Si (2015) and interpolated the velocity vector data at the mesh vertex positions with weights proportional to the vicinity to four closest velocity vector data points for each mesh vertex respectively. The mesh vertices acted as ground truth to train and evaluate our model.
@@
Black blood MRI scans were performed using the following parameters: TR = 10 ms, TE = 3.4 ms, flip angle = 6° and 0.5 mm $\times$ 0.5 mm $\times$ 0.5 mm spatial resolution. The k-space was under-sampled R = 5 times with a Poisson disk pattern and reconstructed using a compressed sensing algorithm for which the previously developed PROUD scanner software patch was used Peper et al. (2020); Gottwald et al. (2020). Reconstruction was carried out in MATLAB, using MRecon (Gyrotools, Zurich, Switzerland) and the BART reconstruction toolbox. In all 3D black blood MRI volumes, the internal carotid artery (ICA), external carotid artery (ECA) and common carotid artery (CCA) were automatically segmented using a previously developed deep learning-based algorithm Alblas et al. (2021). Meshes were subsequently cut at 3 cm below and 2 cm above the carotid bifurcation to match the region-of-interest obtained from 4D flow MRI.
@@
The final dataset consisted of 128 left and 154 right carotid arteries, for which paired meshes were available from 4D flow MRI and black blood MRI. For both 4D flow and black blood MRI, we obtained centerline points by binning and averaging geodesic segments of the artery surface, computed via the vector heat method Crane et al. (2017). Connecting these centerline points by line elements allows us to query the distance to the centerline and thus vessel radius using trigonometry.
@@ 3.2 Machine learning model
In the following section we introduce our machine learning model for learning to predict velocity fields in 3D artery models. We propose a neural network that maps a collection of input features, defined on a set of points, which describe (local) artery geometry within the artery lumen, to a collection of output features representing 3D velocity vectors of blood flow. Note that neither rotation nor translation of the input geometry should influence the relative structure of the predicted vector field. We address this by a steerable, $\mathrm{E}(3)$-equivariant neural network that processes descriptive, geometric features independent of orientation in ambient space. Furthermore, we include the Navier-Stokes equations as loss regularisation.
@@ 3.2.1 Input features
We describe arterial geometry locally by point-wise feature vectors $f^{p}\in\mathbb{R}^{c}$ for each point $p\in\mathcal{P}_{0}$, where $c$ is the channel size and $\mathcal{P}_{0}$ a finite set of points within the artery lumen. We describe each point by its relative position (compare Suk et al. (2023)) to the artery inlet, lumen wall, outlets and centerline. We additionally encode relative position to the ICA outlet, ECA outlet and append a field of zeros (or ones) if the artery is a left (or right) carotid artery (respectively). We do so to account for possible physiological differences with respect to upstream or downstream vasculature and their effect on the flow field. All these features represent relative positions and are thus invariant under translation of the 3D geometry.
@@
Furthermore, we include boundary condition features, as constant 1D scalar fields extended over all points. In particular, we use the norm of the mean velocity vector over the artery inlet and the norm of the standard deviation vector over the artery inlet. This is to simulate velocity measurements in an axial slice of the artery. Such measurements are commonly used in CFD to get subject-specific flow boundary conditions, often via ultrasound Lopes et al. (2020). Conditioning on different inflow velocities is important because boundary conditions vary between subjects and greatly influence blood flow.
@@ 3.2.2 Architecture
We propose an encoder-decoder model (following PointNet++ Qi et al. (2017)) with a contracting and expanding pathway. Given an un-ordered set $\mathcal{P}_{0}=\{p^{1},\dots,p^{n}\}$ of point coordinates $p=(p_{1},p_{2},p_{3})^{\mathsf{T}}\in\mathbb{R}^{3}$ within a subject’s artery lumen, we build a nested hierarchy of $h$ sub-sampled sets
@@ 3.2.3 Steerable equivariant PointNet++
We introduce an equivariant extension of PointNet++ using steerable representations. Generally, a function is called equivariant to a symmetry group, if group actions applied to the inputs result in the equivalent transformation of the outputs. By choosing $\phi$, $\psi$ as steerable MLPs we, in turn, render our PointNet++ $\mathrm{O}(3)$-equivariant. Steerable MLPs that are equivariant under transformations of the orthogonal group $\mathrm{O}(3)$ (rotations and reflections) can be constructed by interleaving Clebsch-Gordan tensor products, gated activation functions and batch normalisation Brandstetter et al. (2022). They require expressing feature vectors throughout the network as steerable tensors (e.g. collections of scalars $s$ and 3D vectors $v$)